OpenStudy (spyro517):

If so then here is the next problem. Line AB contains points A (−1, 3) and B (5, 3). Line AB has a slope that is zero undefined positive negative

OpenStudy (anonymous):

ZERO

OpenStudy (spyro517):

Thank you.

OpenStudy (mathstudent55):

The slope of the line through points \((x_1, y_1)\) and \((x_2, y_2) \) is \(slope = m = \dfrac{y_2 - y_1}{x_2 - x_1} \) In your case, your points are A(-1, 3) and B(5, 3), so the slope is \(slope = m = \dfrac{3 - 3}{5 - (-1)} = \dfrac{0}{6} = 0 \)